One number is twice as large as the second. The sum of the two numbers is 366. What is the larger number?

Answer by mananth(16946)

let one number be x
the other number will be 2x
x+2x=366
3x=366
/3
x=122
the larger number is 2 *122=244
..
m.ananth@hotmail.ca